OpenStudy (anonymous):

Amy sold pens and pencils for a fundraiser at school. Pencils sold for $1.50 and pens sold for $2. She sold a total of 35 items and made $61.50. How many pens did Amy sell? 16 17 18 19

OpenStudy (anonymous):

17

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

I think you have to form a let statement

OpenStudy (anonymous):

wait its 18

OpenStudy (anonymous):

ill teel you how

OpenStudy (anonymous):

take 18 and muliplty it by the cost of the pens you get 36 take 61.5(the total) and subtract 36 from it and you get 25.5 now take 25.5 and divide it by the cost of the pencils which is $1.50 and you get 17 which is the number of pencils sold

OpenStudy (anonymous):

I thought u said it was 18 now its 17 again?

OpenStudy (anonymous):

no thats how many pencils he sold

OpenStudy (anonymous):

he sold 18 pens and 17 pencils take the cost of the 18 pens ($36)m and then add that to the amount of 17 pencils($25.5) and you would get your grand total of $61.5

OpenStudy (anonymous):

does it make sense now?

OpenStudy (anonymous):

let p = pens let c = pencils p + c = 35 1.5c + 2p = 61.50 p = 35 - c 1.5c + 2(35-c) = 61.50 -0.5c = -8.5 c = 17 So the amount of pencils sold is 17 thus the amount of pens sold is 18.

OpenStudy (anonymous):

there is another way to put it